# Inferring Pluggable Types with Machine Learning

### v1

@@ III Encodings
This section details how we encoded the input programs for the two kinds of models we consider: the graph-based models via our novel type-system-specific NaP-AST encoding (which is described via the methodology to construct a NaP-AST encoding for an arbitrary type system in section III-A) and the text-based models via prompting (section III-B).
III-A Graph-based Encoding: the NaP-AST
@@
This section describes our novel methodology for building a type-system-specific program encoding for the graph models (GCN and GTN), which we call a Name-Augmented, Pruned Abstract Syntax Tree (NaP-AST). The encoding only includes information relevant to the given type system, but is in a form suitable for training a deep learning model. The overall methodology for NaP-AST construction is shown in fig. 1.
@@
Figure 1: Order of steps in NaP-AST construction: 1. Extract AST (section III-A1). 2. Prune irrelevant nodes or types via proof (section III-A2) and via ablation studies. (section III-A3) 3. Augment the pruned AST by connecting the variables with identical names (section III-A4). 4. Prune irrelevant statements on the basis of relevance (section III-A5). 5. Cluster the ASTs by graph structure (section III-A6). 6/7. Train or predict (section III-A7). 8. If predicting, post-process to ensure consistency (section III-A8).
The base of a NaP-AST is the program’s Abstract Syntax Tree (AST). The first step (section III-A1) to construct a NaP-AST is to augment the AST with additional nodes that model information traditionally stored in the AST directly (for example, whether a declaration is public or private), thereby making that information available to the deep learning model later. This is necessary because the internal complexity of the AST data structure produced by JavaParser is not visible to the deep learning model.
Figure 2: In heterogeneous graphs (i.e. having multiple edge types), a meta-path is a sequence of connected nodes and edges, where each edge has a distinct type.
@@
The next step in constructing a NaP-AST is pruning information that is not relevant to the property of interest from the graph produced by the previous step. We do this by conducting a series of targeted ablation experiments: we iteratively remove particular types of AST nodes (and all associated edges) from the NaP-AST and retrain a preliminary model. Node types whose removal does not negatively impact the F1 score of the preliminary model are removed from all NaP-ASTs. Shrinking the size of the NaP-AST by removing irrelevant information is critical to the scalability of our approach, since the model architectures we consider are highly sensitive to the size of the input graph. There are three pruning phases:
  1. 1.
@@
    prune nodes that are provably unrelated to the property of interest (section III-A2).
  2. 2.
@@
    use an ablation experiment to prune node types that are unrelated to the target property (section III-A3). This pruning step occurs before name augmentation when processing the AST for a particular class.
  3. 3.
@@
    use an ablation experiment on program statement node types to prune entire subtrees under each type of node, as well as related nodes (section III-A5). This pruning step occurs after name augmentation when processing the AST for a particular class, so we defer discussing it until after we introduce the name-augmentation step.
Each pruning phase is theoretically optional, but we found that all three were necessary to produce NaP-AST graphs that are small enough to realistically train a deep learning model when we conducted our feasibility study.
@@
The last part of NaP-AST construction is augmenting the graph with edges that encode program elements that have the same name, which is described in section III-A4. The goal of this step is to encode basic dataflow information in a way that is visible to a deep learning model. For example, this stage enables a model to determine that a parameter checked against null in a method should be annotated as @Nullable by connecting the check to the parameter’s declaration.
@@
Throughout this section, we assume the existence of a training dataset of human-annotated code that contains type qualifiers for the type system of interest. We discuss whether such datasets exist for real type qualifiers in section IV.
III-A1 Abstract Syntax Tree Extraction
@@
For each class in the training dataset, we first generate its Abstract Syntax Tree (AST). Traversing the AST, we generate a graph consisting of nodes and edges. Unlike HiTyper’s Type Dependency Graph [9], we have not attempted to build a graph of type dependencies. We have been able to obtain good results without creating a complex data structure. We think this is because pluggable type systems are much simpler than the type system of a language. Each node is represented by a node id, node type and a special field indicating whether the node contained a type qualifier in the input data (and if, which type qualifier). Any actual AST nodes for the type qualifiers generated by the parser are discarded, so that the human-written type qualifiers are not visible to the model during training. Edges in this graph are parent-child connections represented as an adjacency list.
@@
The graph contains two types of edges: parent-child, and child-parent. Representing the internals of each block or statement by a different edge type prevents relevant properties from leaking out into proximate but otherwise unrelated nodes in the AST. On the other hand, we do want statements using the same variables to be proximate in the graph. We achieve this later using the name layer described in section III-A4.
@@
We then add additional, useful information that is traditionally stored in the AST nodes themselves—in particular, modifiers (e.g., public, static, etc.) —to the graph’s structure, which makes it available to the models during training. We created a single node in the graph for each modifier and connected the modifiers to the relevant nodes using a multi-hot vector encoding. While this may seem like an implementation detail, we mention it here because practical parsers often store this important information not in the AST itself but in fields that are lost if naïvely converted to a graph where each node has one type. We can overcome this issue by representing both the node type and the modifier using multi-hot encoding.
III-A2 Pruning Phase 1: Remove Provably-Unrelated Information
@@
When considering a particular type system, we usually know that some parts of the program are definitely completely irrelevant to the property of interest. If we formalize that notion at the AST level, irrelevant parts of the program can be immediately removed from the AST. For example, a pluggable type system that tracks whether integers are non-negative (which would be useful as e.g., part of a larger system for typechecking array accesses) need not consider parts of the program that do not interact with integers. Therefore, the first pruning stage of NaP-AST construction is to remove any nodes that are provably irrelevant. The set of nodes that are irrelevant will differ when constructing a NaP-AST for each different type system; this pruning stage exists to allow someone following our methodology to build an inference model for a new type system to prune anything they know is irrelevant right away, before the more complex pruning stages.
III-A3 Pruning Phase 2: Ablation by Node Type
@@
Like some Java types are irrelevant for predicting pluggable types, so are some AST node types. For example, MemberValuePair and IntersectionType nodes produced by JavaParser would be irrelevant for predicting non-compound types. To summarize the AST further, we conduct an ablation study on node types. For each node type, we drop it and calculate the F1 score on a held-out test set. We did the same test by dropping random nodes. For node types which get a higher F1 score than dropping random nodes, we drop them from the AST.
III-A4 Name Augmentation
@@
To inform the model of the identity of a candidate node, we layer a graph with different node and edge types on top of the graph derived from the AST in previous steps. This layer connects the names of the variables to nodes that make use of those variables, using different node and edge types. This improves accuracy by allowing the model to connect uses of names to their declarations. Since type annotations need to be placed on declarations, this step is critical: the model must know that a particular test or assignment is related to a specific declaration. For example, fig. 3 shows an example for a variable named "x", which is connected as a separate graph layered over the pruned AST.
The names are connected using a different edge type. The second set of edges are stored in a different adjacency matrix.
@@
MethodDeclarationParameterAssignExpr"x"
Figure 3: Name Augmentation. This diagram shows the pruned AST as yellow nodes, and a node ("x") in the name layer in green. All the nodes that use the same name are connected by a second edge and node type.
III-A5 Pruning Phase 3: Ablation of Statement Subtrees
Even after the pruning phases described above, after name augmentation the resulting graph is prohibitively large to train the graph-based models in section II. So, we need another pruning strategy. Our key insight at this stage is that many statements in the program are irrelevant to the property of interest. For example, parts of the program that do not include any null literals are unlikely to be related to nullability. However, dropping all such statements hurts model performance (as measure by F1 score) too much, so we performed an ablation study to determine which kinds of statements are safe to prune. Our goal at this stage is to prune statements, their corresponding subtrees, and any other related nodes. We use the name-augmentation layer added to the graph in section III-A4 to determine which non-subtree nodes are related to a statement: when we prune a statement and its subtree, we also prune any node that is directly connected to the statement’s subtree via a name layer.
III-A6 Clustering
In most machine learning projects, one cleans the dataset by pruning outliers. Instead of trying to visualize the different kinds of graphs so that we can prune outliers by hand, we clustered the dataset using k-means clustering and trained a separate model for each cluster. We chose a value of k=5 by minimum within-cluster variation with the elbow method [20]. This method raised the overall F1 score by an average of approximately 0.5, so it is critically-important to overall model performance.
III-A7 Prediction
@@
III-A8 Post-processing
@@
Deep learning methods are good for finding recommendations through hidden patterns in the data. Unfortunately, their outputs are not guaranteed to be internally consistent. Post-processing steps remove obvious inconsistencies: they certainly do not guarantee that the model’s outputs will be sensible. They should, however, be conservative in a sense: the annotations they add or remove from the model’s output must be guaranteed to make the resulting annotation set for the whole program more consistent.
III-B Text Encoding
To have Magicoder place annotations, we could not provide it with complete source files, because many were too large. We extracted segments of program text that fit in its 2048 token limit (alongside our prompt) using simple rules requiring the resulting text be syntactically complete. Without these rules, Magicoder will break most code trying to fix syntactic errors. Our prompting code’s rules try to include ending braces and complete loops, methods or files where possible.
For each extracted segment, we gave a simple prompt, "Add \$ANNOannotations to the following Java code snippet where having them wouldimprove code quality. Enclose code in ‘‘‘java...‘‘‘" (where \$ANNO is the literal name of the desired annotation: @Nullable in our feasibility study). If "‘‘‘java...‘‘‘" was in the output, the part of the output contained within was written back into the appropriate location in the program. Otherwise, we wrote back the original unannotated code.
This prompting strategy is relatively primitive. There are two reasons for this: first, the prompt is competing for space in the token limit with the code. A longer prompt requires us to provide less code at a time to the model, reducing the model’s performance. Second, the LLM’s performance was poor enough relative to the graph-based models in our preliminary experiments that it did not seem worthwhile to further optimize it. It can therefore be regarded as a baseline in the experiments in the following sections.
IV Dataset
@@
To evaluate the encodings in section III and to answer our other two research questions—i.e., which model is best suited to the task and how much data is actually required—we performed a feasibility study on the most popular pluggable type system: preventing null-pointer dereferences in Java using a @Nullable annotation. Multiple typecheckers for preventing this problem exist and have been deployed in industry, including Uber’s NullAway [2], Meta’s NullSafe [3], and the Checker Framework’s Nullness Checker [4, 5].
@@
The first challenge in predicting nullability annotations is collecting training data. Each nullability checker has its own set of annotations. Further, there are annotations provided by frameworks (e.g., the JSR 305 nullability annotations in the JDK, or those provided by Android) that do not have an associated checking tool. This fragmentation in the ecosystem means that for any single nullability checker or annotation, there may not be sufficient training data to build an effective model. (This mirrors the general situation for pluggable typecheckers, wherein most checkers are used by a relatively small number of projects, leading to a paucity of training data when considering whether training a model is practical.)
@@
Our key insight in dataset preparation is that the various nullability tools all use similar definitions of @Nullable: it universally means “this element may or may not be null.” Since the most widely-used pluggable typecheckers use @NonNull as the default type for unannotated code (e.g., both NullAway [2] and the Checker Framework [4, 5] use this defaulting scheme), bringing a program into compliance with a typechecker usually requires a human to write mostly or only @Nullable annotations. The consequence of this defaulting scheme is that we can restrict a model’s goal to the prediction of only @Nullable annotations: we do not need to directly predict @NonNull or any of the other, more complex annotations used by various checking tools. In turn, the need to only predict @Nullable means that we can combine training data that uses different @Nullable annotations, as long as those annotations have the same semantics. We used every annotation in the comprehensive list of @Nullable annotations with the same semantics in the Checker Framework’s manual [21].
@@
We used the search feature of Sourcegraph’s CLI tool [22]to collect Java classes that import various @Nullable annotations. We searched for @Nullable annotations by themselves. However, this approach led to a dataset where Jetbrains projects were overrepresented. We think this happened because Jetbrains has the most popular projects that use the @Nullable annotation. We switched to searching for a variety of Nullable imports to gather a diverse dataset. We do not require code in our training set be checked by some nullability tool: instead, we try to model where a human would write a nullability annotation. E.g., @Nullable annotations that a human wrote as documentation are also useful as training data.
Some classes, despite importing a nullability annotation, did not actually contain any such annotations. After selecting only those classes that did, we removed very small and very large classes. We removed very small classes because in initial experiments, we found that keeping them in the training set harms model performance: it assumes that all small classes have all @Nullable members. We also removed classes that were too large to fit our training pipeline. After all of these removals, we were left with 32,370 classes; we removed 60,292 classes in total. Each class in this dataset contains one or more @Nullable annotations.
V NaP-ASTs for Nullability

### v2

@@ 3. Encodings
  • our novel type-system-specific NaP-AST encoding (which is described via the methodology to construct a NaP-AST encoding for an arbitrary type system in section 3.2), and
  • the text-based encoding via prompting (section 3.3).
3.1. Type Dependency Graph
Prior work (Peng et al., 2022b) used the TDG encoding for inferring type annotations for programs in dynamically-typed languages such as Python. The TDG is our baseline encoding for graph models: it has worked for a similar problem in the past. A TDG has four node types: 1) symbol nodes, which represent variables for which the types need to be inferred. Each occurrence of a variable is treated as a unique node to handle scenarios where the variable’s type may change at runtime. 2) expression nodes, which represent expressions that generate types, such as arithmetic operations, function calls, and list comprehensions. 3) branch nodes, which represent branching points in data flow, typically corresponding to conditional statements. 4) merge nodes, which represent merging points in data flow, such as the convergence of different branches of execution. Our implementation of TDGs for Java involved specific adjustments to handle Java-specific syntax and semantics. For example, we have removed the possibility that a variable’s types may change at runtime. We trained TDGs on a GCN model. A GTN would not confer much benefit since the edges of a TDG have relatively homogeneous semantics. However, the number of types inferred for a pluggable type system is small in comparison to the use case a TDG was designed for. Therefore, the main functionality proposed by these systems is underutilized for predicting pluggable types, motivating our NaP-AST encoding. An example TDG is shown in fig. 2.
3.2. Our Novel Graph Encoding: the NaP-AST
@@
This section describes our novel methodology for building a type-system-specific program encoding for the graph models (GCN and GTN), which we call a Name-Augmented, Pruned Abstract Syntax Tree (NaP-AST). Unlike a TDG, this encoding only includes information relevant to the property being enforced by a particular pluggable type system, in a form suitable for training a deep-learning model. The overall methodology for NaP-AST construction is shown in fig. 3. We explain how we built a NaP-AST encoding for a nullability type system in detail later, in section 5; this section gives an overview of the process of NaP-AST construction for an arbitrary pluggable type system.
The base of a NaP-AST is the program’s Abstract Syntax Tree (AST). The first step (section 3.2.1) to construct a NaP-AST is to augment the AST with additional nodes that model information traditionally stored in the AST directly (for example, whether a declaration is public or private), thereby making that information available to the deep-learning model later. This is necessary because the internal complexity of the AST data structure produced by JavaParser is not visible to the deep-learning model.
@@
Figure 3. Order of steps in NaP-AST construction.
@@
The next step in constructing a NaP-AST is pruning information that is not relevant to the property of interest from the graph produced by the previous step. We do this by conducting a series of targeted ablation experiments: we iteratively remove particular types of AST nodes (and all associated edges) from the NaP-AST and retrain a preliminary model using a small dataset (976 classes, in our instantiation of this process; this number is related to batch sizes and memory limitations on the training machine).
The F1 score was calculated by comparing the predicted nodes with the ground truth using the values in the confusion matrix, namely the harmonic mean of precision and recall. The F1 score on this preliminary model was used to determine which nodes to prune.
Node types whose removal does not negatively impact the F1 score of the preliminary model are removed from all NaP-ASTs. Shrinking the size of the NaP-AST by removing irrelevant information is critical to the scalability of our approach, since the model architectures we consider are highly sensitive to the size of the input graph. There are three pruning phases:
  1. (1)
@@
    prune nodes that are provably unrelated to the property of interest (section 3.2.2).
  2. (2)
@@
    use an ablation experiment to prune node types that are unrelated to the target property (section 3.2.3). This pruning step occurs before name augmentation when processing the AST for a particular class.
  3. (3)
@@
    use an ablation experiment on program statement node types to prune entire subtrees under each type of node, as well as related nodes (section 3.2.5). This pruning step occurs after name augmentation when processing the AST for a particular class, so we defer discussing it until after we introduce the name-augmentation step.
Each pruning phase is theoretically optional, but we found that all three were necessary to produce NaP-AST graphs that are small enough to realistically train a deep-learning model. If the pruning steps were not applied, less relevant code would be available to the model in each step (because the model is memory-bound). This makes the model’s performance suffer during prediction: in our early experiments on a subset of the full dataset, we found that omitting pruning reduced the model’s F1 score from 0.95 to 0.78.
@@
The last NaP-AST construction step is augmenting the graph with edges encoding program elements with the same name, which is described in section 3.2.4. The goal of this step is to encode basic dataflow information in a way that is visible to a deep-learning model. For example, this stage lets a model determine that a parameter checked against null in a method should be annotated as @Nullable by connecting the check to the parameter’s declaration.
@@
In this section, we assume the existence of a training dataset of human-annotated code that contains type qualifiers for the type system of interest. We discuss whether such datasets exist for real type qualifiers in section 4.
3.2.1. Abstract Syntax Tree Extraction
@@
For each class in the training dataset, we first generate its Abstract Syntax Tree (AST). Traversing the AST, we generate a graph consisting of nodes and edges. Unlike HiTyper’s Type Dependency Graph (Peng et al., 2022a), we have not attempted to build a graph of type dependencies. We have been able to obtain good results without creating a complex data structure. We think this is because pluggable type systems are much simpler than the type system of a language. Each node is represented by a node id, node type and a special field indicating whether the node contained a type qualifier in the input data (and if, which type qualifier). Any actual AST nodes for the type qualifiers generated by the parser are discarded, so that the human-written type qualifiers are not visible to the model during training. Edges in this graph are parent-child connections represented as an adjacency list.
@@
The graph contains two types of edges: parent-child, and child-parent. Representing the internals of each block or statement by a different edge type prevents relevant properties from leaking out into proximate but otherwise unrelated nodes in the AST. On the other hand, we do want statements using the same variables to be proximate in the graph. We achieve this later using the name layer described in section 3.2.4.
@@
We then add additional, useful information that is traditionally stored in the AST nodes themselves—in particular, modifiers (e.g., public, static, etc.) —to the graph’s structure, which makes it available to the models during training. While this may seem like an implementation detail, we mention it here because practical parsers often store this important information not in the AST itself but in fields that remain unrepresented in the graph structure. We represent this information by creating a single node in the graph for each modifier and connecting the modifiers to the relevant nodes using a multi-hot vector encoding.
3.2.2. Pruning Phase 1: Remove Provably-Unrelated Information
@@
When considering a particular type system, we usually know that some parts of the program are definitely completely irrelevant to the property of interest. If we formalize that notion at the AST level, irrelevant parts of the program can be immediately removed from the AST. For example, a pluggable type system that tracks whether integers are non-negative (which would be useful as e.g., part of a larger system for typechecking array accesses) need not consider parts of the program that do not interact with integers. Therefore, the first pruning stage of NaP-AST construction is to remove any nodes that are provably irrelevant. For example, primitive types cannot store null values in Java, and so NullGTN removes them at this stage. The set of nodes that are irrelevant will differ when constructing a NaP-AST for each different type system; this pruning stage exists to allow someone following our methodology to build an inference model for a new type system to prune anything they know is irrelevant right away, before the more complex pruning stages.
3.2.3. Pruning Phase 2: Ablation by Node Type
@@
Like some Java types are irrelevant for predicting pluggable types, so are some AST node types. For example, MemberValuePair and IntersectionType nodes are irrelevant for predicting non-compound types. To summarize the AST further, we conduct an ablation study on node types. For each node type, we drop it and calculate the F1 score on a held-out test set. We did the same test by dropping random nodes. We drop node types which score worse than dropping random nodes from the AST. Note that when we prune a node, we connect the node’s children to its parents.
3.2.4. Name Augmentation
@@
To inform the model of the identity of a candidate node, we layer a graph with different node and edge types on top of the graph derived from the AST in previous steps. This layer connects the names of the variables to nodes that make use of those variables, using different node and edge types. This improves accuracy by allowing the model to connect uses of names to their declarations. Since type annotations need to be placed on declarations, this step is critical: the model must know that a particular test or assignment is related to a specific declaration. For example, fig. 4 shows an example for a variable named ”x”, which is connected as a separate graph layered over the pruned AST. The names are connected using a different edge type. The second set of edges are stored in a different adjacency matrix.
@@
3.2.5. Pruning Phase 3: Ablation of Statement Subtrees
Even after the pruning phases described above, after name augmentation the resulting graph is prohibitively large to train the graph-based models in section 2. So, we need another pruning strategy. Our key insight at this stage is that many statements in the program are irrelevant to the property of interest. For example, parts of the program that do not include any null literals are unlikely to be related to nullability. However, dropping all such statements hurts model performance (as measure by F1 score) too much, so we performed an ablation study to determine which kinds of statements are safe to prune. Our goal at this stage is to prune statements, their corresponding subtrees, and any other related nodes. We use the name-augmentation layer added to the graph in section 3.2.4 to determine which non-subtree nodes are related to a statement: when we prune a statement and its subtree, we also prune any node that is directly connected to the statement’s subtree via a name layer.
3.2.6. Clustering
In machine learning projects, one typically cleans the dataset by pruning outliers. Instead of trying to visualize the different kinds of graphs so that we can prune outliers by hand, we clustered the dataset using k-means clustering and trained a separate model for each cluster. We chose a value of k=5 by minimum within-cluster variation with the elbow method (Cui et al., 2020). This method raised the overall F1 score by approximately 0.5, so it is critically-important to overall model performance.
3.2.7. Prediction
@@
MethodDeclarationParameterAssignExpr”x”
Figure 4. Name Augmentation. This diagram shows the pruned AST as yellow nodes, and a node (”x”) in the name layer in green. All the nodes that use the same name are connected by a second edge and node type.
3.2.8. Post-processing
@@
Deep-learning methods are good for finding recommendations through hidden patterns in the data. Unfortunately, their outputs are not guaranteed to be internally consistent. Post-processing steps remove obvious inconsistencies: they certainly do not guarantee that the model’s outputs will be sensible. They should, however, be conservative in a sense: the annotations they add or remove from the model’s output must be guaranteed to make the resulting annotation set for the whole program more consistent.
3.2.9. Other types
The NaP-AST encoding can be adapted with relevant changes to any type system with finite types. For example, consider a type system for the signedness of integers. The relevant node types, etc., would almost certainly differ from the things that matter for null safety (e.g., null literals might not matter, but addition or subtraction nodes might matter a lot), so the process that we describe in section 3 would need to be repeated. We separated the description of the process in section 3 from the instantiation of that process for null safety in section 5 to facilitate repeating our approach in the future.
3.3. Text Encoding
To have LLMs place annotations, we could not provide it with complete source files, because many were too large. We extracted segments that fit within a conservative  2 K-token budget to keep inference consistent across models and reduce nonsensical outputs, even though some models (e.g., CodeQwen 1.5-7B-Chat) support much larger contexts (up to 64 K).
To avoid hallucinations that change code structure, we restrict the output of LLMs to yes or no, using them as classifiers. First, the script extracts potential candidates for @Nullable annotation from the Java file by scanning its lines with two regular expressions, field declarations and method signatures. The latter captures both return types and parameter lists. For each match, a unique identifier and their corresponding elements are stored in a dictionary for further processing. Next, a prompt is constructed consisting of a concise system message (”You are a helpful assistant.”) followed by a detailed user message. The user message provides the full Java source code and lists each extracted candidate element, prefixed by its unique key. The LLM is then asked to decide for each candidate whether it should be annotated with @Nullable. The script explicitly instructs the LLM to return a strictly formatted JSON object that maps each candidate key to either ”yes” or ”no.” For any candidate marked with ”yes,” it inserts the @Nullable annotation in the appropriate location, whether at a field, method return, or parameter. Without these guardrails, we found that the LLMs break most code segments trying to make syntactic changes—for example, by adding or removing import statements unnecessarily. Our prompting code’s rules try to include ending braces and complete loops, methods or files where possible to help reduce the impact of these problems.
4. Dataset
@@
To evaluate the encodings in section 3 and to answer our other two research questions—i.e., which model is best suited to the task and how much data is actually required—we performed a feasibility study on the most popular pluggable type system: preventing null-pointer dereferences in Java using a @Nullable annotation. Multiple typecheckers for preventing this problem exist and have been deployed in industry, including Uber’s NullAway (Banerjee et al., 2019b), Meta’s NullSafe (Pianykh et al., 2022), and the Checker Framework’s Nullness Checker (Papi et al., 2008; Dietl et al., 2011).
@@
The first challenge in predicting nullability annotations is collecting training data. Each nullability checker has its own set of annotations. Further, there are annotations provided by frameworks (e.g., the JSR 305 nullability annotations in the JDK, or those from Android) without an associated checking tool. This fragmentation in the ecosystem means that for any single nullability checker or annotation, there is not enough training data to build an effective model. This mirrors the general situation for pluggable typecheckers, wherein most checkers are used by only a small number of projects, leading to too little training data when considering whether training a model is practical.
@@
Our key insight in dataset preparation is that the various nullability tools all use similar definitions of @Nullable: it universally means “this element may or may not be null.” Since the most widely-used pluggable typecheckers use @NonNull as the default type for unannotated code (e.g., both NullAway (Banerjee et al., 2019b) and the Checker Framework (Papi et al., 2008; Dietl et al., 2011) use this defaulting scheme), bringing a program into compliance with a typechecker usually requires a human to write mostly or only @Nullable annotations. Consequently, we can restrict a model’s goal to the prediction of only @Nullable annotations: we do not need to directly predict @NonNull or any other, more complex annotations. In turn, the need to only predict @Nullable means that we can combine training data that uses different @Nullable annotations, as long as those annotations have the same semantics. We used every annotation in the comprehensive list of @Nullable annotations with the same semantics in the Checker Framework’s manual (cf-, 2024).
@@
We used the search feature of Sourcegraph’s CLI tool (Slack and Liu, 2013) to collect Java classes that import @Nullable annotations. We searched for @Nullable annotations by themselves. However, this approach led to a dataset where JetBrains projects were overrepresented. We think this happened because Jetbrains has the most popular projects that use the @Nullable annotation. We switched to searching for a variety of Nullable imports to gather a diverse dataset. We do not require code in our training set be checked by some nullability tool: instead, we try to model where a human would write a nullability annotation. E.g., @Nullable annotations that a human wrote as documentation are also useful as training data.
Some classes, despite importing a nullability annotation, did not actually contain any such annotations. After selecting only those classes that did, we removed very small and very large classes. We define “very small” classes as those with fewer than 300 AST nodes. There are 11,019 such classes in the dataset. Based on manual review, we observed that these classes lacked nontrivial logic (besides getters, setters and annotated fields). Our early experiments showed that retaining them in the training set skews the model towards annotating every annotatable location. We define “very large” classes as those with more than 8,000 nodes after pruning. We decided to remove these because we wanted the model to see the complete code of each class during training. 8,000 nodes was the empirical limit imposed by the memory capacity of the training machine. After all of these removals, we were left with 32,370 classes; we removed 60,292 classes in total. Note that despite excluding both categories during training, the models operate on all classes regardless of size when deployed. The NullAway benchmark (section 7.1) contains 1641 “too small” classes and 24 “too large” classes that would have been removed during training. Our results there show that removing these classes during training does not cause problems during deployment. Each class in this dataset contains one or more @Nullable annotations. There are 217,922 @Nullable annotations in these classes in total.
5. NaP-ASTs for Nullability